OpenStudy (anonymous):

what are the 3 possible points you can use to graph y= 3-x

OpenStudy (anonymous):

HELP!!!!

jimthompson5910 (jim_thompson5910):

You will have infinitely many points because you can plug in any number for x to get a paired corresponding value for y

jimthompson5910 (jim_thompson5910):

for instance, say x = 0 y = 3 - x y = 3 - 0 ... plug in x = 0 y = 3 so the point (0,3) is on the line

jimthompson5910 (jim_thompson5910):

now say x = 1 y = 3 - x y = 3 - 1 ... plug in x = 1 y = 2 now we know that the point (1,2) is on the line

jimthompson5910 (jim_thompson5910):

What happens when x = 3?